solve:
2y^2+xy=5
4y+x=7

Answer by stanbon(75887)

solve:
2y^2+xy=5
4y+x=7
-----------------------
Solve the 2nd equation for "x":
x = 7-4y
----------
Substitute that into the 1st equation and solve for "y":
2y^2+(7-4y)y = 5
2y^2+7y-4y^2 = 5
2y^2-7y+5 = 0
y = [7 +- sqrt(49-4*2*5)]/4
y = [7 +- 3]/4
y = 5/2 or y = 1
----------------------
No solve for "x":
If y = 5/2, x = 7-4*(5/2) = -3
If y = 1, x = 7-4*1 = 3
----------------------------
Solutions: (-3,5/2) or (3,1)
=========================
Cheers,
Stan H.